# An Survey of the Evolution of Language Model-Based Dialogue Systems

### v1

@@ 1 Introduction
  • Early stage – Statistical language models (SLMs). The first dialogue system — Eliza, was proposed by MIT in 1966, which was even earlier than the rise of SLMs in the 1990s, allowing plausible conversations between humans and machines. Subsequently, a series of dialogue systems were constructed, including a travel agent aiding the client to make a simple return trip to a single city in California [15] and a rule-based chatbot configured with a persona, employing heuristic pattern matching rules to address inquiries about ages, hobbies, interests, and etc [16]  [^3] https://web.njit.edu/~ronkowit/eliza.html. In addition to academic endeavors, more and more companies pay attention to the field of dialogue systems and have developed their own products. For example, IBM has developed WATSON, Apple has Siri, and Microsoft has XiaoIce. These early virtual assistants have limited functionalities and follow a rigid process, primarily assisting users with single and simple tasks like scheduling appointments, setting reminders, and providing basic answers to questions. At this stage, most of these dialogue systems are task-oriented, the prevailing nature of these dialogue systems is predominantly task-oriented, employing either modular-based approaches or machine-learning methods based on SLMs.
  • The independent development of TOD and ODDNeural language models (NLMs). Around 2015, there was a sea change when the sequence to sequence (seq2seq) framework was proposed, in which salient features were learned jointly with the training of the model — NLM [17, 18]. Specifically, NLMs characterize the probability of word sequences by neural networks, e.g., two typical recurrent neural networks (RNNs) such as long short-term memory networks (LSTM) [19] and gated recurrent units (GRU) [20]. These NLMs serve as the backbone to map variable-length input sequences to variable-length output sequences which is achieved by employing two main components: an encoder and a decoder. This paradigm shift facilitated the emergence of Open-domain Dialogue (ODD) systems ($\sim$2015, Figure  2), garnering attention due to a well-defined end-to-end framework enabled by the seq2seq network. Simultaneously, NLMs were also applied to TOD, with the goal of achieving better performance and performing complex tasks in multiple domains [21, 22, 23].
  • Fusions of Dialogue Systemspre-trained language models (PLMs). Traditional NLMs requires extensive hand-crafted features selection and domain-specific knowledge, making them cumbersome and time-consuming to develop for each task. However, with the advent of deep learning and large-scale language corpora, pre-trained language models (PLMs) emerged as a groundbreaking solution. These models are pre-trained on massive amounts of unlabeled text data, capturing rich semantic and syntactic patterns [24, 11, 25]. By leveraging pre-training, subsequent fine-tuning on specific tasks becomes more accessible, enabling the models to adapt quickly and achieve impressive results with minimal task-specific data. In the realm of dialogue systems, DialoGPT exemplifies this approach by undergoing pre-training on 147 million conversation-like exchanges extracted from Reddit comment chains and then finetuning on a few training examples to generate more relevant, informative, and context-consistent responses [13]. This paradigm, commonly known as pre-train and finetune [26], leads to the resulting model as pre-trained dialogue model (PDM) that undergoes fine-tuning using a dialogue corpus. It is worth noting that the major differences between PDMs lies in the the corpus used for training and the backbone language models adopted, such as PLATO [27, 28], BlenderBot [29], Meena [30], Pangu-bot [31], PLUG [32], and etc. In addition, there is an increasing focus on the fusion of different tasks inside TOD (e.g., End-2-end TOD), TOD with ODD, and propose Unified Dialogue System (UniDS) in this period (Figure 2).
  • LLM-based dialogue systemLarge language models (LLM). Recognizing the correlation between increased pre-training corpus size and model size with enhanced performance across various NLP tasks [33], researchers try to scale the model size and the size of pre-training corpus simultaneously to be more sample efficient, enabling the model to learn more intricate patterns and representations from the data. Thus, the PLMs become Large Language Models (LLMs), such as GLM [34], LLaMA [35], and InstructGPT [4]. Due to the larger model size and the utilization of extensice, high-quality pre-training corpora, these LLMs offer unprecedented capabilities in many language understanding and generation tasks, including but not limited to question-answering and named entity recognition, thus reshaping the landscape of dialogue systems. With further instruction-tuning using data as (instruction, input, output), these LLMs are capable of answering diverse questions and following various instructions such as writing an email and telling a joke. At this moment, the LLMs can be directly used as the dialogue systems. Moreover, finetuning the LLM with in-domain data, i.e., dialogue/conversational corpus, can further enhance its conversational ability. For example, the evolution of GPT-3.5 to InstructGPT, and then ChatGPT exactly follows this path. Similarly, ChatGLM [36] is also finetuned using publicly available Chinese dialogue datasets based on the corresponding LLM – GLM. To enhance alignment with human preferences and values, the incorporation of reinforcement learning becomes imperative, particularly in the final stages of model refinement. This involves leveraging feedback from either human evaluators [4] or automated systems [37] to iteratively improve the model’s performance and ensure its adherence to desired criteria. After all, this type of LLM-based dialogue system (LLM-based DS) performs well at both task-oriented dialogue and chit chat, making it an ideal foundation for a general conversational AI.
It is worth noting that the development of LM-based Dialogue System is a ongoing and evolving process, and the stages are not rigidly separated by specific turning points. Instead, there are overlaps and continuous advancements across these stages, particularly in the later stages. Generally, with the advancement of LM, the boundary between TOD and ODD and the boundary between DM and LM have become increasingly blurred, ushering in a new era of LLM-based Dialogue System. Despite that the existing studies have thoroughly examined the development and challenges associated with different types of dialogue systems, such as TOD [2] and ODD [1], there still remains a notable gap in the literature when it comes to offering a holistic perspective on the evolutionary trajectory of dialogue systems, particularly in light of the advancements in LM. By critically analyzing and understanding the impact of pivotal milestones, we can gain valuable insights into the possibilities that await and the transformative impact dialogue systems may have on diverse domains and applications in the foreseeable future. This survey aims to present a comprehensive overview of the realm of LM-based dialogue systems  [^4] We have made thorough efforts to cover a wide range of works, but acknowledge the possibility of omissions. Readers are encouraged to provide suggestions for any missed contributions or errors. We are committed to updating the article with new approaches or definitions proposed by the community., shedding light on the prominent directions which the field is currently progressing. This entails a meticulous review of the existing literature, aimed at elucidating the paradigm shifts observed within different stages of LM-based dialogue systems. These shifts encompass a range of integrations across various levels, notably encompassing the task, data, and framework dimensions.
Figure 2: The number of publication at ACL Anthology conferences of different research directions in LM-based dialogue systems (The raw data is downloaded around 31/10/2023).
The remainder of this survey is organized as follows: Section 2 provides the description of different types of dialogues and define the problem in a unified manner, followed by the early development of two types of LM-based dialogue systems: TOD and ODD in Section 3. Sections  4 review and summarize the evolution of different dialogue systems when PLMs are scaled to LLMs. We then investigate emerging trends and open challenges in LLM-based DS (See Section 5). Then, Section 6 discusses the open problems for LLM-based dialogue systems. Finally, we conclude the survey in Section 7 by summarizing the major findings and discussing the remaining issues for future work.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hiddendraw, align=left, minimum width=2.5em, s sep=6pt, inner xsep=2pt, inner ysep=1pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=2em,font=, where level=2text width=6.5em,font=, where level=3text width=16em,font=, where level=4text width=16em,font=, [LM-based
@@
Understanding (NLU) [RNN: RNN-LU [38, 39]
LSTM: BiLSTM [40, 41], Attention-BiLSTM [42]
Slot-Gated [43] , DF-Net [44] [BERT: Joint-NLU, [45] BERT-Joint [46]
StackNLU [47]
RoBERTa: MultiDomain [48]
Others: MinDS [49] ] ] ] [Dialogue State
Tracking (DST) [RNN: Word-level RNN [50], NBT [51], MD-DST
[52], MemN2N [53], GLAD [54]
GRU: TRADE [55], SAS [56] ,
LSTM: FJST [57], [Bert: CSFN [58], SUMBT [59], EMD [60],
DS-DST [61]
Albert: DiCoS [62]
GPT-2: NP-DST [63], SimpleToD [9],
SOLOIST [64] ] ] ] [Dialogue Policy
Learning (DPL) [RNN: LRU [65], RS-RNN [66], MADPL [67]
@@
LSTM: OnlineRL [68], GDPL [69], AdvL [70] [BERT: PPO-OFF-Comb [37]
GPT2: SimpleToD [9] ] ] ] [Natural Language
@@
Generation (NLG) [ LSTM: SC-LSTM [71], Meta-NLG [72], HDSA [73] [GPT2: SimpleToD [9], SC-GPT [74]] ] ] [End-to-End (E2E) [LSTM: Network [22] [See Section 4.1] ] ] ] [ODD [ Generative DS [ RNN: NRM [75], Ext-ED [76]
LSTM: NCM [77], P-NCM [78], PersonaChat [79]
GRU: CoPerHED [80], STAR-BTM [81],
CKG-aware [82], K-NCM [83]
] ] [ Retrieval-based DS [ RNN: SMN [95]
LSTM: DL2P [96], Match-LSTM [97],
Attentive-LSTM [98]
GRU: Multi-view [99], DUA [100],
CNN: ARC [101], RSTP [102], LSM [103]
] ] [ Hybrid [ RNN: XiaoIce [112], EnsembleGAN [113]
GRU: Ensemble [114], PTE [115]
LSTM: EED [116], RR [117], HybridNCM [118]

### v2

@@ 1. Introduction
2. Background
Before we proceed, it is essential to figure out how many different types of dialogs are in our daily life. Table 1 demonstrates some examples of four major types of dialogs, including:
  1. (a)
    social chit-chat. This type of dialog involves casual and informal conversations aimed at establishing and maintaining social connections. It includes light-hearted discussions about topics such as current events (Sharma and et al., 2020; Sun and et al., 2021), personal experiences (Zhang et al., 2018; Xu et al., 2022a), hobbies, and interests (Shang et al., 2015). Social chit-chat is typically used to create a friendly atmosphere and build social bonds, requiring the DS to generate persona-consistent, empathetic, and compassionate responses besides helpful.
  2. (b)
    knowledge-grounded chit-chat. Knowledge-grounded dialogs revolve around exchanging information and discussing topics that require a certain level of expertise or specific knowledge. It can involve discussing academic subjects, professional domains, or specialized areas of interest. This type of dialog often focuses on sharing and acquiring knowledge (Dinan et al., 2019; Zhou and et al., 2020). The knowledge can be formed in different formats such as graphs (Bosselut et al., 2019), tables (Cheng and et al., 2022), figures (Antol and et al., 2015), or even a combination of them (Wu and et al., 2021). Most of previous methods for this type of dialog based on retrieval-augmented generation (RAG) framework (Lewis and et al., 2021).
  3. (c)
    question answering. Question-answering dialogs involve a user asking specific questions, and the dialogue system providing relevant and accurate answers. These dialogs are typically focused on obtaining specific information or clarifications. The system’s role is to understand the user’s queries and provide appropriate responses based on the available knowledge or resources (Sun and et al., 2021; Lin et al., 2022).
  4. (d)
    task-oriented dialogue. Task-oriented dialogues center around assisting users in accomplishing specific tasks or goals. Users seek help from the dialogue system to perform actions such as making reservations, booking flights, ordering products, or seeking recommendations. These dialogues often involve a series of steps or sub-tasks that need to be completed to achieve the overall objective (Peng and et al., 2017; Budzianowski and et al., 2018; Zhu et al., 2020).
These are general descriptions of the different types of dialogue, and there may be variations or combinations of these types in real-world dialogue scenarios (Liu et al., 2022b). The diversity and complexity of each type of dialogue not only contribute to the wide range of applications for dialogue systems but also give rise to various challenges and problems. These challenges pose unique requirements and considerations, necessitating the development of dialogue systems that are both versatile and robust that can effectively handle the intricacies of each dialogue scenario in a unified manner.
2.1. Task Definition
Following the unified definition recently (Liu et al., 2022b; Zhang and et al., 2023a), we formulate the task of dialogue system as a seq2seq problem by directly mapping the dialogue context to the final response with or without the assistance of different external sources such as the local database in TOD, different knowledge bases in ODD, e.g., persona and documents. Formally, given the dialogue context $c=\{u_{1},s_{1},...,u_{t}\}$, and optionally the external knowledge bases $K$, the goal of the dialogue system is to generate a helpful and harmless response $r$ as shown below:
(1) $r=\boldsymbol{M}(c,K)\quad or\quad\boldsymbol{M}(c)$
Where $\boldsymbol{M}$ is parameterized by different language models. Specifically, there are significant differences in interactions between various dialogue systems with the external knowledge base $K$. In this work, we discuss two major interactions corresponding to TOD and ODD respectively. Formally, the TOD first generates the belief state according to the given context (Zhao et al., 2022; Peng et al., 2021):
(2) $b=\boldsymbol{M}(c)$
(3) $TOD(r|c,K)=f(r|c,b,s)f(b|c)$
Secondly, for the ODD, the processing is relatively simpler by directly rewriting the dialogue context as the query or using the whole dialogue context to search the external knowledge base $K$ (Dinan et al., 2019). In this case, TOD can be regarded as one special type of ODD with the source of knowledge as corresponding database (Tian et al., 2022).
(4) $ODD(r|c,K)=f(r|c,q,s)f(q|c)$
It is important to clarify here that ODD contains the cases which does not require any external knowledge, by directly mapping dialogue context to corresponding responses using LMs (Chen and et al., 2017).
2.2. Metrics and Benchmarks
To evaluate the performance of different types of dialogue systems, several metrics and benchmarks are carefully designed. In order to assess various aspects of dialogue generation, including task success, response quality, and consistency. Depending on the type of dialogue system, the evaluation metrics and benchmarks may differ significantly. Table 2 shows the common metrics and benchmarks used in the dialogue systems.
Table 2. Metrics and benchmarks for different tasks in dialogue systems.
DS Task Benchmarks Metrics Evaluation Type TOD NLU SNIPS (Coucke and et al., 2018), ATIS, Few-NERD (Ding and et al., 2021) Acc, F1 Automatic DST DSTC-{2,4} (Henderson et al., 2014a; Kim and et al., 2017), Wizard-of-Oz (Mrkšić et al., 2017), CoSQL (Yu and et al., 2019) Joint-Acc Automatic DPL MultiWoZ (Eric and et al., 2020), RisaWoZ (Quan and et al., [n. d.]), CrossWoZ (Zhu et al., 2020) Acc, F1, Reward Automatic NLG E2E (Novikova and et al., 2017), BAGEL (Mairesse and et al., 2010), RNNLG (Wen and et al., 2016), FewshotWoZ (Peng et al., 2020) BLEU, Rouge, METEOR Automatic General MultiWoZ (Eric and et al., 2020), RisaWoZ (Quan and et al., [n. d.]), CrossWoZ (Zhu et al., 2020), SGD (Rastogi and et al., 2020) Succ, Turns Automatic KddRES (Wang and et al., 2021a), CamRest676 (Wen et al., 2017a), Frames (El Asri and et al., 2017), KvRET (Eric and Manning, 2017) Multilingual: BiToD (Lin and et al., 2021a), X-RisaWoZ (Moradshahi and et al., 2023), IndoToD (kau, 2023) ODD Retrieval WoW (Dinan et al., 2019), FoCus (Jang and et al., 2022), KdConv (Zhou and et al., 2020), KBP (Wang and et al., 2023c) Recall@k, Hit@k Automatic NLG Chit-chat: DailyDialog (Li and et al., 2017), Topical-Chat (Gopalakrishnan and et al., 2023), CMU_DoG (Zhou et al., 2018) Empathy: EmpatheticDialogues (Rashkin and et al., 2019), ESConv (Liu and et al., 2021), D4 (Yao et al., 2022) Persona: PersonaChat (Zhang et al., 2018), KvPI (Song and et al., 2020), PersonalDialog (Zheng et al., 2019), FoCus (Jang and et al., 2022) BLEU, Rouge, PPL, Distinct Automatic Memory: DuLeMon (Xu et al., 2022a), MSC (Xu et al., 2022b), PerLTQA (Du and et al., 2024) Fluency, Coherence Human Evaluation Knowledge: WoW (Dinan et al., 2019), KdConv (Zhou and et al., 2020), OpenDialKG (Moon and et al., 2019), KBP (Wang and et al., 2023c) Win Rate LLM-as-a-Judge Multilingual: XPersona (Lin and et al., 2021b), XDailyDialog (Liu and et al., 2023c) Multimodal: Image-Chat (Shuster and et al., 2020), MMChat (Zheng and et al., 2022), StickerConv (Zhang and et al, 2024) NLI PersonaChat (Zhang et al., 2018), KBP (Wang and et al., 2023c), KvPI (Song and et al., 2020) Consistency Automatic Dialogue NLI (Welleck and et al., 2019), FoCus (Jang and et al., 2022) LLM-as-a-Judge
2.2.1. Task-oriented Dialogue System.
A modular task-oriented dialogue system consists of several sub-tasks, each responsible for a specific function. Similarly, an end-to-end task-oriented dialogue system performs these functions implicitly, but the evaluation metrics used to assess each module in a modular system can also be applied to evaluate the corresponding aspects of an end-to-end system.
Natural Language Understanding (NLU)
NLU processes user utterances and extracts relevant semantic frames, consisting of two subtasks: intent classification and slot filling. NLU can be rule-based, where predefined patterns or rules are used to extract information, or it can employ deep learning techniques like intent classification and named entity recognition. The result usually includes the intent of user and corresponding slots information such as departure location and destination if the user’s intent is detected as book flight ticket.
  • Metrics. To evaluate the performance for these two tasks (i.e., intent detection and slot filling), Accuracy and F1-score are widely employed as defined as follows:
    (5) $Acc=\frac{\sum_{i=1}^{N}\mathbf{1}(\hat{y}_{i}=y_{i})}{N}\quad\text{F1}=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$
    where $\hat{y}_{i}$ is the predicted label for the $i_{th}$ instance, $y_{i}$ is the ground truth label, $N$ is the total number of instances and $\mathbf{1}$ is the indicator function.
  • Benchmarks. SNIPS (Coucke and et al., 2018) and ATIS (atis) are two classic and commonly used datasets for intent detection and slot filling tasks in NLU, and there are some other datasets such as Few-NERD (Ding and et al., 2021) which focus on few-shot named entity recognition.
Dialogue State Tracking (DST)
DST maintains and updates the state of the conversation as it progresses. It keeps track of the user’s goals, preferences, and any relevant context obtained from the dialogue context by updating the intent and slots information. DST can use a rule-based approach to update according to the results from the NLU module, or replace the NLU module and directly track the dialogue state from a dialogue context with similar architecture as NLU (Balaraman and et al., 2021; Wang et al., 2021).
  • Metrics. It is crucial to track the dialogue states in terms of joint accuracy of the intent and slots as the conversation progresses. Joint accuracy measures the correctness of the dialogue state at each turn, considering it correct only if all intents and slot values are predicted accurately (Henderson et al., 2014a).
  • Benchmarks. The Dialogue System Technology Challenges (DSTC) have introduced several well-known benchmarks like DSTC2 (Henderson et al., 2014a) and DSTC4 (Kim and et al., 2017), primarily focusing on dialogue state tracking. In addition to these, there are more advanced benchmarks designed to tackle complex scenarios, including multi-domain dialogue state tracking (Wizard-of-Oz (Mrkšić et al., 2017)) and tasks grounded in SQL for more sophisticated state management (CoSQL (Yu and et al., 2019)).
Dialogue Policy Learning (DPL)
The dialogue policy component determines the system’s actions or responses based on the current dialogue state. It maps the dialogue state to appropriate system actions, which can include asking clarification questions, providing recommendations, or executing relevant API calls. Dialogue policies can be rule-based, handcrafted, or learned using techniques like reinforcement learning (Peng and et al., 2017; Takanobu and et al., 2019, 2020; Wang et al., 2020b, 2022) or supervised learning (Henderson and et al., 2008; Shang et al., 2015). More details can be found in the survey paper (Kwan and et al., 2023).
  • Metrics. Besides the above mentioned Accuracy and F1, DPL additionally uses reward score as one of evaluation metrics to evaluate whether or not current policy leads to greater reward scores.
  • Benchmarks. It is common to directly use some general task-oriented dialogue corpus to evaluate the performance of dialogue policy, including but not limited to MultiWoZ (Eric and et al., 2020), RisaWoZ (Quan and et al., [n. d.]) and CrossWoZ (Zhu et al., 2020).
Natural Language Generation (NLG)
NLG generates system responses in natural language based on the dialogue policy’s output. It transforms structured information or system actions into coherent and fluent sentences that can be understood by users. NLG can employ template-based approaches, rule-based generation, or more advanced techniques like neural network-based generation models (Li et al., 2020; Zhu and et al., 2020a).
  • Metrics. Ultimately, these generated response can be directly compared with the ground truth response in the original dataset. The quality is commonly measured using BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and sometimes METEOR (Banerjee and et al., 2005).
  • Benchmarks. The commonly used benchmarks for natural language generation (NLG) in task-oriented dialogue systems include E2E NLG (Novikova and et al., 2017), BAGEL (Mairesse and et al., 2010), RNNLG (Wen and et al., 2016), FewshotWoZ (Peng et al., 2020). Among these, E2E NLG and BAGEL focus on single-domain scenarios, while RNNLG and FewshotWoZ are designed for multi-domain settings.
General
In addition to independently evaluating the performance of each sub-task in a modular task-oriented dialogue system, there are several general metrics and benchmarks available for assessing the overall performance of the entire system. To ensure a comprehensive and consistent evaluation, ConvLab (Zhu and et al., 2020a) has been introduced as an open-source toolkit. It provides researchers with the capability to build task-oriented dialogue systems and evaluate the performance of each component effectively.
  • Metrics. The primary objective of a task-oriented dialogue system is to assist users in completing predefined tasks, making the success rate the most critical metric, defined as the proportion of dialogues in which the system successfully helps users achieve their goals. Another important metric is the average number of dialogue turns, as an effective system should complete tasks in as few turns as possible, minimizing unnecessary interactions.
  • Benchmarks. There are different types of task-oriented dialogue corpus in different applications: i) general domain: MultiWoZ (Eric and et al., 2020), RisaWoZ (Quan and et al., [n. d.]), CrossWoZ (Zhu et al., 2020); and ii) specific domain: KddRES (Wang and et al., 2021a), CamRest676 (Wen et al., 2017a), Frames (El Asri and et al., 2017), KvRET (Eric and Manning, 2017); and iii) multilingual: BiToD (Lin and et al., 2021a), X-RisaWoZ (Moradshahi and et al., 2023), IndoToD (kau, 2023). Specifically, KddRES (Wang and et al., 2021a) is the first Cantonese task-oriented dialogue corpus on restaurant, while BiToD (Lin and et al., 2021a) is the first bilingual multi-domain dataset for end-to-end task-oriented dialogue modeling. All of these benchmarks can be used for evaluation of specific sub-tasks or general end-to-end performance.
2.2.2. Open-domain Dialogue System
Open-domain dialogue systems (ODDs) are designed to engage in conversations on a wide range of topics without domain-specific constraints. Evaluating the performance of these systems is challenging due to the complexity and variability of dialogue responses. To ensure comprehensive assessment, evaluation typically involves three key tasks: Knowledge Retrieval, Natural Language Generation (NLG), and Natural Language Inference (NLI). Each task has dedicated metrics and benchmarks that measure system performance from different perspectives. While different types of ODDs may only involve one or two of these sub-tasks, we showcase the common process typically integrates all three to ensure accurate and comprehensive evaluation.
Knowledge Retrieval.
Knowledge retrieval is a fundamental step in ODDs, particularly for knowledge-grounded dialogue systems that rely on external sources to generate informed and contextually relevant responses. The goal is to retrieve the most relevant knowledge from a large-scale knowledge base or document collection.
  • Metrics. Retrieval tasks typically use Recall@k and Hit@k, measuring how often the correct candidate response appears in the top-k retrieved items. For example, Recall@k is:
[…]
  • Benchmarks. Many knowledge-augmented dialogue corpora include additional annotations to support retrieval tasks. Examples include WoW (Dinan et al., 2019), FoCus (Jang and et al., 2022), KdConv (Zhou and et al., 2020), and KBP (Wang and et al., 2023c). These datasets often provide relevant documents, sentences, or knowledge graphs to facilitate effective information retrieval and dialogue generation.
Natural Language Generation (NLG)
After optionally retrieving relevant knowledge, the system can generate a response. Since open-ended dialogues allow for multiple valid responses, it is unreasonable to only consider single ground truth answer. This makes evaluating response quality challenging due to the diversity of plausible responses.
  • Metrics. There are three different ways to evaluate the generated response for open-domain dialogue system. First of all, there are several automatic evaluation metrics, including BLEU, Rouge, PPL and Distinct-N. Secondly, since the content of dialogue is open-ended, there is no strict ground truth response given the dialogue context. Thus, it is usually required to conduct additional human evaluation to directly employ human annotators to grade the quality of generated responses based on different aspects, such as fluency and coherence. Recently, as LLMs demonstrate a high level of agreement with human judgments, they are increasingly being used to replace human annotators in grading responses (i.e., LLM-as-a-judge). For example, these LLMs can be employed to directly compare responses generated by different systems, ultimately determining the winning model based on win rate.
  • Benchmarks. There are numerous benchmarks designed to evaluate the performance of open-domain dialogue systems. These benchmarks can be categorized based on whether they rely on external knowledge and the specific types of knowledge sources they use. Some datasets, such as those for chit-chat (DailyDialog (Li and et al., 2017), Topical-Chat (Gopalakrishnan and et al., 2023),) and empathetic dialogues (EmpatheticDialogues (Rashkin and et al., 2019), ESConv (Liu and et al., 2021), D4 (Yao et al., 2022)), do not require external knowledge, focusing instead on generating engaging and emotionally aware conversations. In contrast, other datasets are designed to incorporate various types of external knowledge into responses, including world knowledge (WoW (Dinan et al., 2019), KdConv (Zhou and et al., 2020), KBP (Wang and et al., 2023c)), persona (PersonaChat (Zhang et al., 2018), KvPI (Song and et al., 2020), PersonalDialog (Zheng et al., 2019)), and memory (DuLeMon (Xu et al., 2022a), MSC (Xu et al., 2022b), PerLTQA (Du and et al., 2024)). Additionally, several benchmarks explore more complex scenarios, such as multilingual conversations (XPersona (Lin and et al., 2021b) and XDailyDialog (Liu and et al., 2023c)) and multi-modal interactions (MMChat (Zheng and et al., 2022)).
Natural Language Inference (NLI)
Sometimes, it is required to assess whether the generated response is consistent with the retrieved knowledge. A dedicated NLI model is often used to classify the relationship between the response and the knowledge as consistent, contradictory, or neutral.
  • Metrics. The consistency score can be calculated as the proportion of responses classified as consistent out of all generated responses. Depending on specific type of incorporated knowledge, there are different types of consistency scores such as persona consistency, knowledge consistency and memory consistency.
  • Benchmarks. It is worthy noting that mostly the NLI benchmark can be extracted from existing open-domain dialogue corpus if it contains the augmentation of external knowledge, such as KBP (Wang and et al., 2023c) and FoCus (Jang and et al., 2022). In addition, there are several general dialogue NLI datasets such as Dialogue NLI (Welleck and et al., 2019).
3. The Independent Development of TOD & ODD
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hiddendraw, align=left, minimum width=2.5em, s sep=6pt, inner xsep=2pt, inner ysep=1pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=2em,font=, where level=2text width=6em,font=, where level=3text width=14em,font=, where level=4text width=12em,font=, [LM-based
@@
Understanding (NLU) [RNN: RNN-LU (Yao et al., 2013; Mesnil and et al., 2014)
LSTM: BiLSTM (Yao et al., 2014; Hakkani-Tür and et al., 2016), Attention-BiLSTM (Liu and Lane, 2016)
Slot-Gated (Goo et al., 2018) , DF-Net (Qin et al., 2020) [BERT: Joint-NLU, (Chen et al., 2019) BERT-Joint (Castellucci and et al., 2019)
StackNLU (Qin and et al., 2019)
RoBERTa: MultiDomain (Qin and et al., 2022)
Others: MinDS (Gerz et al., 2021) ] ] ] [Dialogue State
Tracking (DST) [RNN: Word-level RNN (Henderson et al., 2014b), NBT (Mrkšić et al., 2017), MD-DST
(Mrkšić et al., 2015), MemN2N (Perez and Liu, 2017), GLAD (Zhong et al., 2018)
GRU: TRADE (Wu et al., 2019a), SAS (Hu et al., 2020) ,
LSTM: FJST (Eric and et al., 2020), [BERT: CSFN (Zhu and et al., 2020b), SUMBT (Lee et al., 2019), EMD (Wang et al., 2021),
DS-DST (Zhang and et al., 2020a)
Albert: DiCoS (Guo and et al., 2022)
GPT-2: NP-DST (Ham et al., 2020), SimpleToD (Hosseini-Asl et al., 2020),
SOLOIST (Peng et al., 2021) ] ] ] [Dialogue Policy
Learning (DPL) [RNN: LRU (Su and et al., 2015a), RS-RNN (Su and et al., 2015b), MADPL (Takanobu and et al., 2020)
@@
LSTM: OnlineRL (Su et al., 2016), GDPL (Takanobu and et al., 2019), AdvL (Liu and Lane, 2018) [BERT: PPO-OFF-Comb (Wang et al., 2022)
GPT2: SimpleToD (Hosseini-Asl et al., 2020) ] ] ] [Natural Language
@@
Generation (NLG) [ LSTM: SC-LSTM (Wen and et al., 2015), Meta-NLG (Mi et al., 2019), HDSA (Chen and et al., 2019) [GPT2: SimpleToD (Hosseini-Asl et al., 2020), SC-GPT (Peng et al., 2020)] ] ] [End-to-End (E2E) [LSTM: Network (Wen et al., 2017b) [See Section 4.1] ] ] ] [ODD [ Generative DS [ RNN: NRM (Shang et al., 2015), Ext-ED (Parthasarathi and Pineau, 2018)
LSTM: NCM (Vinyals and Le, 2015), P-NCM (Li et al., 2016), PersonaChat (Zhang et al., 2018)
GRU: CoPerHED (Kottur et al., 2017), STAR-BTM (Zhang et al., 2020a),
CKG-aware (Zhong et al., 2021), K-NCM (Ghazvininejad et al., 2018)
] ] [ Retrieval-based DS [ RNN: SMN (Wu and et al., 2017)
LSTM: DL2P (Yan and et al., 2016), Match-LSTM (Wan et al., 2016),
Attentive-LSTM (Tan et al., 2016)
GRU: Multi-view (Zhou et al., 2016), DUA (Zhang and et al., 2018),
CNN: ARC (Hu et al., 2014), RSTP (Severyn and Moschitti, 2015), LSM (Shen et al., 2014)
] ] [ Hybrid [ RNN: XiaoIce (Zhou et al., 2020), EnsembleGAN (Zhang and et al., 2019)
GRU: Ensemble (Song and et al., 2018), PTE (Wu et al., 2019b)
LSTM: EED (Pandey et al., 2018), RR (Weston et al., 2018), HybridNCM (Yang and et al., 2019)
] ] ] [FuseDS [ TOD $\rightarrow$ ODD [ UniDS (Zhao et al., 2022), HyKnow (Gao and et al., 2021), ACCENTOR (Sun et al., 2021) FusedChat (Young et al., 2022), text width=25em ] ] [ ODD $\rightarrow$ TOD [ Q-ToD (Tian et al., 2022), DuClarifyDial (Liu et al., 2022b), GODEL (Peng et al., 2022), DialogStudio (Zhang and et al., 2023a), MDS (Nehring and et al., 2021), text width=25em ] ] ] ]